# SpaCE-10: A Comprehensive Benchmark for Multimodal Large Language Models in Compositional Spatial Intelligence

### v2

@@ 1 Introduction
3DQA fe-3dgqa SCN scannet - 902 Sim. $\times$× $\times$× $\times$×
ScanQA scanqa SCN scannet 167 10k Sim. $\checkmark$ $\times$× $\times$×
FE-3DGQA fe-3dgqa SCN scannet 100 3.9k Sim. $\times$× $\times$× $\times$×
SQA3D sqa3d SCN scannet 132 6.9k Sim. $\checkmark$ $\times$× $\times$×
CLEVER3D clever3d SCN scannet 133 10k Sim. $\times$× $\times$× $\times$×
3D-LLM 3d-llm OBJ,SCN scannet ,HM3D hm3d - 30k Sim. $\times$× $\times$× $\times$×
M3DBench m3dbench SCN scannet - 1.5k Sim.+LLM $\times$× $\times$× $\times$×
MSQA msqa SCN scannet ,3RS 3rscan ,ARK arkitscenes 381 3.5k LLM $\times$× $\times$× $\times$×
VSI vsi SCN scannet ,3RS 3rscan ,SCN++ scannet++ 288 5.0k Acc. $\times$× $\times$× $\checkmark$
SpaCE-10 (Ours) SCN scannet ,3RS 3rscan ,ARK arkitscenes ,SCN++ scannet++ 811 6k Acc. $\checkmark$ $\checkmark$ $\checkmark$
@@ 3.2 Hierarchical Annotation Pipeline
[…]
@@ 4.2 Overall Results
Human vs. MLLMs. We first test human performance on SpaCE-10, taking the average score of two humans to obtain the ‘Human’ score in Tab. 2. The results indicate that although human performance does not meet expectations when facing with more complex numerical and reasoning tasks, the total score of 81.5% is still significantly higher than all existing MLLMs. In comparison, the best open-source MLLM only achieves an average score of 52.8%, and the best close-source models only achieve a score of 49.0%. These results demonstrate that the compositional spatial intelligence of MLLMs is far below the human level.
@@
GPT4Scene-Qwen2-VL-7B gpt4scene 16 30.9 37.7 38.0 38.9 41.6 29.5 28.0 32.5 31.9
@@
GPT-4o-2024-11-20 gpt4 4 58.3 32.8 56.2 58.3 56.2 41.6 52.2 23.7 49.0
Gemini-2.0-Flash-Thinking-Exp gemini 10 34.3 25.6 53.1 42.6 53.8 42.2 46.7 31.2 42.2
Claude-3.7-Sonnet-20250219 claude3 7 46.0 44.3 49.1 46.0 49.1 44.3 49.3 25.0 46.2
@@
InternVL2.5-1B internvl2.5 13 33.0 54.1 18.8 43.6 29.9 30.9 41.0 23.7 36.2
InternVL2.5-2B internvl2.5 18 32.2 26.8 27.0 36.6 28.8 34.0 48.2 36.2 33.8
@@
Qwen2.5-VL-7B-Instruct qwen2.5 19 32.7 36.9 36.9 35.3 32.3 27.6 34.2 27.5 33.3
@@
LLaVA-OneVision-7B llava-onevision 8 37.4 33.8 46.4 57.3 34.5 43.3 61.6 21.2 45.2
@@
InternVL2.5-26B internvl2.5 6 34.3 29.3 62.6 55.4 33.0 50.2 61.8 33.7 47.5
Qwen2.5-VL-32B-Instruct qwen2.5 21 19.9 26.5 48.9 36.8 32.3 31.1 30.1 32.5 32.6
InternVL2.5-38B internvl2.5 5 38.1 36.1 64.4 54.3 36.8 45.6 63.0 37.5 48.7
@@
Qwen2.5-VL-72B-Instruct qwen2.5 12 32.4 34.9 55.7 40.9 32.1 36.5 38.0 33.7 37.9
InternVL2.5-78B internvl2.5 3 27.8 45.0 62.4 64.4 40.3 36.6 67.3 40.0 49.7
@@
Open-Source vs. Close-Source. In close-souce MLLMs experiments, GPT-4o-2024-11-20 gpt4 achieves the best overall performance, ranking 4th with a score of 49.0%. It excels in perception tasks, such as in SA with 56.2%, OO with 58.3%, and OS with 56.2%, indicating strong accuracy in recognizing size and position. Additionally, GPT-4o gpt4 achieves the highest score in EQ among all tested models, with a score of 58.3%. However, it struggles with SP, where it scores the lowest among all tasks, suggesting a limitation in scene-level planning ability. Gemini-2.0-Flash-Thinking-exp gemini and Claude-3.7-Sonnet-20250219 claude3 achieve overall scores of 42.2% and 46.2%, respectively. Compared to these models, Claude-3.7-Sonnet shows a more balanced performance across tasks. In open-source MLLMs, among models with a scale of over 72B, LLaVA-OneVision-72B delivers outstanding performance, ranking first overall with a score of 52.8%. It outperforms other models in nearly all tasks, solidifying its top position. Furthermore, InternVL2.5-78B achieves the second-highest score of 49.7%, showing the increasing competitive edge of open-source models in spatial reasoning. These results suggest that the gap between open-source and close-source models has significantly narrowed, and open-source MLLMs even outperform close-source models, especially in compositional spatial intelligence.
@@
InternVL2.5-1B internvl2.5 18.8 43.6 29.9 30.9 41.0 4.4 3.9 4.7 0.5 11.4 32.8 3.0 0.09
InternVL2.5-2B internvl2.5 27.0 36.6 28.8 34.0 48.0 4.7 2.0 1.3 1.5 3.5 34.9 4.1 0.12
InternVL2.5-8B internvl2.5 50.0 55.0 33.6 41.1 59.1 8.4 14.8 8.7 12.9 1.5 47.8 10.6 0.22
InternVL2.5-38B internvl2.5 64.4 54.3 36.8 45.6 63.0 47.8 37.9 7.3 21.3 46.8 52.8 32.2 0.61
InternVL2.5-78B internvl2.5 62.4 64.4 40.3 36.6 67.3 38.4 33.5 12.1 12.9 45.8 54.2 28.5 0.53
@@
EP $\times$× C1,C2,C8 66.6 75.8 63.4 65.3 67.8
$\checkmark$ +C7,C9 30.9 ($\downarrow$ 35.7) 41.1 ($\downarrow$ 34.7) 25.5 ($\downarrow$ 37.9) 27.6 ($\downarrow$ 37.7) 31.3 ($\downarrow$ 36.5)
FR $\times$× C1,C2,C3,C6,C8 70.9 89.7 85.6 85.7 83.0
@@ 4.3 Capability Analysis
Atomic Capability vs. Compositional Capabilities. Firstly, the Tab. 4 shows that on the 5 types of questions with more compositional abilities, the performance of the four models all drops significantly. For the 3 questions with integrating C7 capability, the models’ accuracy in SA, OO, and OS tasks decreases by 19.1%, 7.6%, and 21.0%, respectively. Despite the drop, the accuracy of SA and OO still remains at a decent level. Secondly, in the EP task with both C7 and C9 abilities integrated, the models’ overall accuracy plummets from 67.8% without integration to 31.3%, a huge drop of 36.5%. Similarly, for the FR task, which with the most compositional capacities, the four models’ average score crashes from 83.0% to 42.8%, an even larger drop of 42.9%.
@@
Spatial Capability Breakdown. To better understand the strengths and weaknesses of the current MLLMs across various atomic capabilities, we constructed a model accuracy and atomic capability score matrix to associate QA accuracy with capabilities. As shown in Fig. 4, from the results, we observe several findings. Overall, different models show similar trends in capabilities such as C1 (Object Recognition), C2 (Spatial Localization), C4 (Size Comparison), C6 (Functional Understanding), and C9 (Reverse Reasoning), with scores in these areas predominantly blue, indicating stable performance across models. Capabilities like C5 (Counting) show green scores, indicating weaker performance across models. Notably, the distribution of capabilities across models is relatively consistent, suggesting that while these capabilities are not fully orthogonal, models have a similar overall grasp of different abilities.
@@
Based on the macro analysis in the first step, we can now focus on the scores of each model. Similar to the accuracy trend, LLaVA-OneVision-72B demonstrates the most powerful performance, achieving the highest average scores in over seven atomic capabilities. Among all its abilities, the one with the smallest gap to human performance is C4, representing 67.9% of LLaVA-OneVision-72B with only a 12.3% difference from humans. For most models, C4 is the best mastered atomic capability, indicating that current models have relatively strong capabilities in size comparison. However, in C5, counting abilities, all models perform poorly, with the highest score reaching only 45.5% of GPT-4o, a 25.7% gap from humans. Therefore, MLLMs’ counting ability needs significant improvement.
Additionally, in terms of average scores, the best model, LLaVA-OneVision-72B, only achieved 52.1%, with a 29.1% gap from humans. The best closed-source model, GPT-4o, even has a 34.2% gap from humans. These results indicate that there is still a large room for improvement in both individual atomic spatial capabilities and integrated spatial intelligence in current MLLMs. This capability analysis provides valuable insights for guiding future enhancements of MLLMs. More experiments are attached in the supplementary materials.
@@ B.1 More Mulitple-Choice Results
InternVL2.5-1B internvl2.5 18.8 43.6 29.9 30.9 41.0 4.4 3.9 4.7 0.5 11.4 32.8 3.0 0.09
InternVL2.5-2B internvl2.5 27.0 36.6 28.8 34.0 48.2 1.3 2.0 4.7 1.5 3.5 34.9 4.1 0.12
@@
InternVL2.5-8B internvl2.5 50.0 55.0 33.6 41.1 59.1 8.4 14.8 8.7 12.9 1.5 47.8 10.6 0.2
InternVL2.5-26B internvl2.5 62.6 55.4 33.0 50.2 61.8 30.5 32.0 13.3 17.3 32.8 52.6 25.2 0.49

### v3

@@ 1 Introduction
3DQA fe-3dgqa SCN scannet - 902 Sim. $\times$ $\times$ $\times$
ScanQA scanqa SCN scannet 167 10k Sim. $\checkmark$ $\times$ $\times$
FE-3DGQA fe-3dgqa SCN scannet 100 3.9k Sim. $\times$ $\times$ $\times$
SQA3D sqa3d SCN scannet 132 6.9k Sim. $\checkmark$ $\times$ $\times$
CLEVER3D clever3d SCN scannet 133 10k Sim. $\times$ $\times$ $\times$
3D-LLM 3d-llm OBJ,SCN scannet ,HM3D hm3d - 30k Sim. $\times$ $\times$ $\times$
M3DBench m3dbench SCN scannet - 1.5k Sim.+LLM $\times$ $\times$ $\times$
MSQA msqa SCN scannet ,3RS 3rscan ,ARK arkitscenes 381 3.5k LLM $\times$ $\times$ $\times$
VSI vsi SCN scannet ,3RS 3rscan ,SCN++ scannet++ 288 5.0k Acc. $\times$ $\times$ $\checkmark$
SpaCE-10 (Ours) SCN scannet ,3RS 3rscan ,ARK arkitscenes ,SCN++ scannet++ 811 5.0k Acc. $\checkmark$ $\checkmark$ $\checkmark$
@@ 3.2 Hierarchical Annotation Pipeline
Basic QA Generation. As we mentioned earlier, the questions in SpaCE-10 are composed of multiple atomic capabilities. However, such highly integrated questions are difficult for current MLLMs to generate directly. Therefore, we propose to first generate a basic version of QA, namely basic QA, and then enhance its embedded capabilities. Next, we will briefly introduce the QA types in SpaCE-10. As shown in Fig. 2, SpaCE-10 consists of 8 QA types, including Entity Quantification (EQ), Region Quantification (RQ), Size Assessment (SA), Object-Object Spatial Relationship (OO), Object-Scene Spatial Relationship (OS), Functional Reasoning (FR), and Spatial Planning (SP). All QA pairs are split into Reasoning and Perception. For the Fig. 2 (Left), perception includes 5 sub-types, accounting for $\sim$70% of the questions, and reasoning includes 3 sub-types, comprising $\sim$30% of the questions.
@@ 4.2 Overall Results
Human vs. MLLMs. We first test human performance on SpaCE-10, taking the average score of two humans to obtain the ‘Human’ score in Tab. 2. The results indicate that although human performance does not meet expectations when facing with more complex numerical and reasoning tasks, the total score of 85.3% is still significantly higher than all existing MLLMs. In comparison, the best open-source MLLM only achieves an average score of 52.8%, and the best close-source models only achieve a score of 49.0%. These results demonstrate that the compositional spatial intelligence of MLLMs is far below the human level.
@@
GPT4Scene-Qwen2-VL-7B gpt4scene 16 30.9 37.7 38.0 38.9 41.6 29.5 28.0 32.5 34.5
@@
GPT-4o-2024-11-20 gpt4 3 58.3 32.8 56.2 58.3 56.2 41.6 52.2 23.7 49.0
Gemini-2.0-Flash-Thinking-Exp gemini 9 34.3 25.6 53.1 42.6 53.8 42.2 46.7 31.2 42.2
Claude-3.7-Sonnet-20250219 claude3 5 46.0 44.3 49.1 46.0 49.1 44.3 49.3 25.0 46.2
@@
InternVL2.5-1B internvl2.5 14 33.0 54.1 18.8 43.6 29.9 26.7 41.0 23.7 35.3
InternVL2.5-2B internvl2.5 21 32.2 26.8 27.0 36.6 28.8 21.7 48.2 36.2 31.4
@@
Qwen2.5-VL-7B-Instruct qwen2.5 18 32.7 36.9 36.9 35.3 32.3 27.6 34.2 27.5 33.3
@@
LLaVA-OneVision-7B llava-onevision 6 37.4 33.8 46.4 57.3 34.5 43.3 61.6 21.2 45.2
@@
InternVL2.5-26B internvl2.5 8 34.3 29.3 62.6 55.4 33.0 29.2 61.8 33.7 43.3
Qwen2.5-VL-32B-Instruct qwen2.5 20 19.9 26.5 48.9 36.8 32.3 31.1 30.1 32.5 32.6
InternVL2.5-38B internvl2.5 7 38.1 36.1 64.4 54.3 36.8 27.4 63.0 37.5 45.1
@@
Qwen2.5-VL-72B-Instruct qwen2.5 11 32.4 34.9 55.7 40.9 32.1 36.5 38.0 33.7 37.9
InternVL2.5-78B internvl2.5 4 27.8 45.0 62.4 64.4 40.3 23.7 67.3 40.0 47.1
@@
Open-Source vs. Close-Source. In close-souce MLLMs experiments, GPT-4o-2024-11-20 gpt4 achieves the best overall performance, ranking 3rd with a score of 49.0%. It excels in perception tasks, such as in SA with 56.2%, OO with 58.3%, and OS with 56.2%, indicating strong accuracy in recognizing size and position. Additionally, GPT-4o gpt4 achieves the highest score in EQ among all tested models, with a score of 58.3%. However, it struggles with SP, where it scores the lowest among all tasks, suggesting a limitation in scene-level planning ability. Gemini-2.0-Flash-Thinking-exp gemini and Claude-3.7-Sonnet-20250219 claude3 achieve overall scores of 42.2% and 46.2%, respectively. Compared to these models, Claude-3.7-Sonnet shows a more balanced performance across tasks. In open-source MLLMs, among models with a scale of over 72B, LLaVA-OneVision-72B delivers outstanding performance, ranking first overall with a score of 52.8%. It outperforms other models in nearly all tasks, solidifying its top position. Furthermore, InternVL2.5-78B achieves the second-highest score of 47.1%, showing the increasing competitive edge of open-source models in spatial reasoning. These results suggest that the gap between open-source and close-source models has significantly narrowed, and open-source MLLMs even outperform close-source models, especially in compositional spatial intelligence.
@@
InternVL2.5-1B internvl2.5 18.8 43.6 29.9 26.8 41.0 4.4 3.9 4.7 0.5 11.4 32.3 3.0 0.09
InternVL2.5-2B internvl2.5 27.0 36.6 28.8 21.7 48.0 4.7 2.0 1.3 1.5 3.5 32.4 4.1 0.13
InternVL2.5-8B internvl2.5 50.0 55.0 33.6 27.1 59.1 8.4 14.8 8.7 12.9 1.5 45.0 10.6 0.24
InternVL2.5-38B internvl2.5 64.4 54.3 36.8 27.4 63.0 47.8 37.9 7.3 21.3 46.8 49.2 32.2 0.65
InternVL2.5-78B internvl2.5 62.4 64.4 40.3 23.7 67.3 38.4 33.5 12.1 12.9 45.8 51.6 28.5 0.55
@@
EP $\times$ C1,C2,C8 66.6 75.8 63.4 65.3 67.8
$\checkmark$ +C7,C9 26.8 ($\downarrow$ 39.8) 27.1 ($\downarrow$ 48.7) 25.5 ($\downarrow$ 37.9) 27.6 ($\downarrow$ 37.7) 26.8 ($\downarrow$ 41.0)
FR $\times$ C1,C2,C3,C6,C8 70.9 89.7 85.6 85.7 83.0
@@ 4.3 Capability Analysis
Atomic Capability vs. Compositional Capabilities. Firstly, the Tab. 4 shows that on the 5 types of questions with more compositional abilities, the performance of the four models all drops significantly. For the 3 questions with integrating C7 capability, the models’ accuracy in SA, OO, and OS tasks decreases by 19.1%, 7.6%, and 21.0%, respectively. Despite the drop, the accuracy of SA and OO still remains at a decent level. Secondly, in the EP task with both C7 and C9 abilities integrated, the models’ overall accuracy plummets from 67.8% without integration to 26.8%, a huge drop of 41.0%. Similarly, for the FR task, which with the most compositional capacities, the four models’ average score crashes from 83.0% to 42.8%, an even larger drop of 42.9%.
@@
Spatial Capability Breakdown. To better understand the strengths and weaknesses of the current MLLMs across various atomic capabilities, we constructed a model accuracy and atomic capability score matrix to associate QA accuracy with capabilities. As shown in Fig. 4, from the results, we observe several findings. Overall, different models show similar trends in capabilities such as C1 (Object Recognition), C2 (Spatial Localization), C4 (Size Comparison), C6 (Functional Knowledge), and C9 (Reverse Reasoning), with scores in these areas predominantly blue, indicating stable performance across models. Capabilities like C5 (Counting) show green scores, indicating weaker performance across models. Notably, the distribution of capabilities across models is relatively consistent, suggesting that while these capabilities are not fully orthogonal, models have a similar overall grasp of different abilities.
@@
Based on the macro analysis in the first step, we can now focus on the scores of each model. Similar to the accuracy trend, LLaVA-OneVision-72B demonstrates the most powerful performance, achieving the highest average scores in over seven atomic capabilities. Among all its abilities, the one with the smallest gap to human performance is C4, representing 67.9% of LLaVA-OneVision-72B with a 22.3% difference from humans. For most models, C4 is the best mastered atomic capability, indicating that current models have relatively strong capabilities in size comparison. However, in C5, counting abilities, all models perform poorly, with the highest score reaching only 45.5% of GPT-4o, a 35.4% gap from humans. Therefore, MLLMs’ counting ability needs significant improvement.
Additionally, in terms of average scores, the best model, LLaVA-OneVision-72B, only achieved 52.1%, with a 33.8% gap from humans. The best closed-source model, GPT-4o, even has a 38.9% gap from humans. These results indicate that there is still a large room for improvement in both individual atomic spatial capabilities and integrated spatial intelligence in current MLLMs. This capability analysis provides valuable insights for guiding future enhancements of MLLMs. More experiments are attached in the supplementary materials.
@@ B.1 More Mulitple-Choice Results
InternVL2.5-1B internvl2.5 18.8 43.6 29.9 26.8 41.0 4.4 3.9 4.7 0.5 11.4 32.8 3.0 0.09
InternVL2.5-2B internvl2.5 27.0 36.6 28.8 21.7 48.2 1.3 2.0 4.7 1.5 3.5 34.9 4.1 0.12
@@
InternVL2.5-8B internvl2.5 50.0 55.0 33.6 27.1 59.1 8.4 14.8 8.7 12.9 1.5 45.0 10.6 0.24
InternVL2.5-26B internvl2.5 62.6 55.4 33.0 29.2 61.8 30.5 32.0 13.3 17.3 32.8 48.4 25.2 0.52